# SpectrumFM: A Foundation Model for Intelligent Spectrum Management

### v1

@@ IV-A Experimental Settings
The SpectrumFM is pre-trained on a setup comprising four NVIDIA RTX 4090 GPUs, utilizing PyTorch and DeepSpeed for efficient training. The hyperparameters for SpectrumFM are as follows. The mask ratio $r$ is set to 15%, the number of signal symbols is set to 128, the number of attention heads $H$ is set to 4, the latent dimension $d$ is set to 256, the feedforward dimension $d_{\text{feed}}$ is set to 512, and the number of SpectrumFM encoder layers $L$ is set to 16. The pre-training phase consists of 10 epochs with a batch size of 256 and a learning rate of 0.001. The AdamW optimizer is employed for optimization, and early stopping is utilized to prevent overfitting. During the fine-tuning stage, the same learning rate of 0.001 and the AdamW optimizer are used to further adapt the model to specific downstream tasks.
@@ IV-B Automatic Modulation Classification Task
TABLE II: The Overall Performance in the AMC Task.
@@ IV-B2 Baselines
  • Transformer [35] leverages the self-attention mechanism to extract features from the input signal.
@@
  • CGDNN [41] composes of a shallow convolutional network, a gated recurrent unit, and a deep neural network, for robust automatic modulation recognition.
  • DAE [42] is an LSTM denoising auto-encoder framework that efficiently extracts robust features from noisy radio signals for the classification of communication technologies and modulation schemes.
@@
  • ResNet [44] is a residual network, a widely used deep learning model.
@@
  • GRU2 [45] is a two layers of the GRU classifier model.
@@ IV-B3 Overall Performance
To evaluate the overall performance of our proposed method for the AMC task, we conduct extensive experiments on three datasets, namely RML2016.10A, RML2016.10B, and RML2016.04C. The results are summarized in TABLE II, which presents the precision, recall, and F1-score for various models. The results underscore the superior capability of our model in the AMC task, achieving notable improvements in precision, recall, and F1-score over other baseline models. Specifically, our model attains F1-scores of 66.20% on RML2016.10A, 65.93% on RML2016.10B, and 73.46% on RML2016.04C, outperforming baselines like Transformer and AMC_Net by 2.65-9.61 percentage points. Notably, it demonstrates robust precision (75.29% on RML2016.10A) and recall (73.37% on RML2016.04C), addressing challenges in class imbalance.
@@ IV-B4 The Performance on Various SNR Conditions
(a) The accuracy of models on different SNR conditions on RML2016.10A dataset.
(b) The accuracy of models on different SNR conditions on RML2016.10B dataset.
(c) The accuracy of models on different SNR conditions on RML2016.04C dataset.
Figure 3: The accuracy of models on different SNR conditions in the AMC task.
To evaluate the robustness of our proposed method, its performance is assessed and contrasted with several baseline models across different SNR values, as illustrated in Fig. 3. The findings highlight the exceptional capability of our method across varying SNR conditions on three distinct datasets, namely, RML2016.10A, RML2016.10B, and RML2016.04C. Particularly, on the RML2016.10A dataset (Fig. 3(a)), our method showcases a marked improvement in accuracy at low SNR levels (-10 to 0 dB), significantly outperforming all other models. The trend continues on the RML2016.10B dataset (Fig. 3(b)), where our method maintains its superior performance across the entire SNR range, demonstrating enhanced reliability. Furthermore, the robustness of our method is particularly evident on the RML2016.04C dataset (Fig. 3(c)), our method not only consistently outperforms other models but also shows a particularly significant advantage under challenging low SNR conditions (-10 to -5 dB). These results collectively emphasize the efficacy of our proposed method in achieving high accuracy across diverse SNR conditions, underscoring its superior performance and resilience in various communication environments.
@@ IV-B5 Performance in Few-shot Conditions
The bar chart in Fig. 6(a) illustrates the classification accuracy of our method, MSNet, and AMC_Net under different training set ratios (10%, 20%, 30%, and 100%). The vertical axis represents accuracy, while the horizontal axis denotes the proportion of the training set used for model training. From the results, it is evident that our proposed method consistently outperforms MSNet and AMC_Net across all training set ratios. Notably, a key observation from the results is that as the training data ratio decreases, the accuracy of MSNet and AMC_Net drops significantly, whereas our proposed method remains remarkably stable. Even when trained with only 10% of the dataset, our model achieves an accuracy comparable to MSNet and AMC_Net trained on 100% of the data. The results highlight the strong few-shot learning capability of our method.
@@ IV-C3 Performance Comparison
TABLE III: Comparison of Precision, Recall and F1-score in the WTC Task.
@@
The experimental results presented in TABLE III demonstrate the superior performance of our proposed model compared to existing methods in the WTC task. Specifically, our model achieves the highest precision (0.8270), recall (0.8253), and F1-score (0.8226), significantly outperforming the baseline methods. Notably, compared to well-established architectures such as ResNet and VGG, our model exhibits substantial improvements, particularly in Recall and F1-score, indicating its enhanced ability to correctly classify instances while maintaining robustness against misclassification.
@@ IV-D3 Performance Comparison
Fig. 9(a) illustrates the performance of different spectrum sensing methods at various SNR levels. All methodologies exhibit poor performance when the SNR is between -10 dB and -4 dB. However, within the range of SNR from -10 dB to 0 dB, our proposed method demonstrates a higher accuracy rate. Beyond 0 dB, all methodologies achieve nearly perfect sensing accuracy.
Fig. 9(b) shows the receiver operating characteristic (ROC) curve of various spectrum sensing method at -4 dB SNR level. It highlights that our method achieves an AUC of 0.97 at -4 dB SNR, surpassing both LSTM and ResNet, demonstrating its superior accuracy in signal detection. At the same false alarm rate, our method also exhibits better detection performance, further underscoring its effectiveness in the SS task.
(a) Sensing accuracy of our model and baseline models at various SNR level.
(b) ROC curve of various spectrum sensing method at -4 dB SNR level.
Figure 9: Performance comparison of our model and baseline methods in the SS task.
@@ IV-E Anomaly Detection Task
Figure 10: The data acquisition and SpectrumFM deployment platform.
@@ IV-E2 Baseline
We leverage adversarial autoencoder (AAE) [46] as the baseline model for AD task. The AAE is a generative model that integrates an autoencoder with a discriminator to enforce a specific feature distribution in the latent space. In AAE, the autoencoder consists of an encoder-decoder pair, where the encoder maps input data to a latent representation, and the decoder reconstructs the original input from this representation. Furthermore, a discriminator is employed to guide the encoder in shaping the latent space to follow a predefined distribution, such as a Gaussian or other structured priors.
@@ IV-E3 Performance Comparison
IV-F Discussion

### v2

@@ IV-A Experimental Settings
The SpectrumFM is pre-trained on a setup comprising four NVIDIA RTX 4090 GPUs, utilizing PyTorch and DeepSpeed for efficient training. The hyperparameters for SpectrumFM are as follows. The number of signal symbols is set to 128, the number of attention heads $H$ is set to 4, the latent dimension $d$ is set to 256, the feedforward dimension $d_{\text{feed}}$ is set to 512, and the number of SpectrumFM encoder layers $L$ is set to 16 [gulati2020conformer]. The mask ratio $r$ is set to 15%, the pre-training phase consists of 10 epochs with a batch size of 256 and a learning rate of 0.001 [devlin-etal-2019-bert]. The AdamW optimizer is employed for optimization, and early stopping is utilized to prevent overfitting. During the fine-tuning stage, the same learning rate of 0.001 and the AdamW optimizer are used to further adapt the model to specific downstream tasks.
As shown in Fig. 6, the pre-training loss decreases smoothly without oscillations across different objectives, indicating stable convergence of the training process.
@@ IV-B Automatic Modulation Classification Task
TABLE III: The Overall Performance in the AMC Task.
@@ IV-B2 Baselines
  • Transformer [vaswani2017attention] leverages the self-attention mechanism to extract features from the input signal.
@@
  • CGDNN [cgdnn] composes of a shallow convolutional network, a gated recurrent unit, and a deep neural network, for robust automatic modulation recognition.
  • DAE [dae] is an LSTM denoising auto-encoder framework that efficiently extracts robust features from noisy radio signals for the classification of communication technologies and modulation schemes.
@@
  • ResNet [resnet] is a residual network, a widely used deep learning model.
@@
  • GRU2 [gru2] is a two layers of the GRU classifier model.
@@ IV-B3 Overall Performance
To evaluate the overall performance of our proposed method for the AMC task, we conduct extensive experiments on three datasets, namely RML2016.10A, RML2016.10B, and RML2016.04C. The results are summarized in TABLE III, which presents the precision, recall, and F1-score for various models. The results underscore the superior capability of our model in the AMC task, achieving notable improvements in precision, recall, and F1-score over other baseline models. Specifically, our model attains F1-scores of 66.20% on RML2016.10A, 65.93% on RML2016.10B, and 73.46% on RML2016.04C, outperforming baselines like Transformer and AMC_Net by 2.65-9.61 percentage points. Notably, it demonstrates robust precision (75.29% on RML2016.10A) and recall (73.37% on RML2016.04C), addressing challenges in class imbalance.
@@ IV-B4 The Performance on Various SNR Conditions
Figure 3: Loss curves during the pre-training stage.
Figure 4: The accuracy of models on different SNR conditions on RML2016.10A dataset.
Figure 5: The accuracy of models on different SNR conditions on RML2016.10B dataset.
Figure 6: The accuracy of models on different SNR conditions on RML2016.04C dataset.
To evaluate the robustness of our proposed method, its performance is assessed and contrasted with several baseline models across different SNR values, as illustrated in Fig. 6-6. The findings highlight the exceptional capability of our method across varying SNR conditions on three distinct datasets, namely, RML2016.10A, RML2016.10B, and RML2016.04C. Particularly, on the RML2016.10A dataset (Fig. 6), our method showcases a marked improvement in accuracy at low SNR levels (-10 to 0 dB), significantly outperforming all other models. The trend continues on the RML2016.10B dataset (Fig. 6), where our method maintains its superior performance across the entire SNR range, demonstrating enhanced reliability. Furthermore, the robustness of our method is particularly evident on the RML2016.04C dataset (Fig. 6), our method not only consistently outperforms other models but also shows a particularly significant advantage under challenging low SNR conditions (-10 to -5 dB). These results collectively emphasize the efficacy of our proposed method in achieving high accuracy across diverse SNR conditions, underscoring its superior performance and resilience in various communication environments.
@@ IV-B5 Performance in Few-shot Conditions
The bar chart in Fig. 9(a) illustrates the classification accuracy of our method, MSNet, and AMC_Net under different training set ratios (10%, 20%, 30%, and 100%). The vertical axis represents accuracy, while the horizontal axis denotes the proportion of the training set used for model training. From the results, it is evident that our proposed method consistently outperforms MSNet and AMC_Net across all training set ratios. Notably, a key observation from the results is that as the training data ratio decreases, the accuracy of MSNet and AMC_Net drops significantly, whereas our proposed method remains remarkably stable. Even when trained with only 10% of the dataset, our model achieves an accuracy comparable to MSNet and AMC_Net trained on 100% of the data. The results highlight the strong few-shot learning capability of our method. Fig. 10(a) illustrates the few-shot fine-tuning performance with 100 samples per class on the AMC task using the RML2016.10A dataset under varying SNR conditions. The results show that our method achieves consistently better accuracy than MSNet, Transformer, and AMC_Net across all SNR levels, demonstrating its robustness and good transferability in low-data regimes.
@@ IV-B6 Learning Efficiency
(a) AMC.
(b) WTC.
Figure 10: Few-shot fine-tuning with 100 samples per class on AMC and WTC tasks in different SNR conditions.
@@ IV-C3 Performance Comparison
TABLE IV: Comparison of Precision, Recall and F1-score in the WTC Task.
@@
Figure 13: Sensing accuracy of our model and baseline models at various SNR level.
Figure 14: ROC curve of various spectrum sensing method at -4 dB SNR level.
Figure 15: The data acquisition and SpectrumFM deployment platform.
The experimental results presented in TABLE IV demonstrate the superior performance of our proposed model compared to existing methods in the WTC task. Specifically, our model achieves the highest precision (0.8270), recall (0.8253), and F1-score (0.8226), significantly outperforming the baseline methods. Notably, compared to well-established architectures such as ResNet and VGG, our model exhibits substantial improvements, particularly in Recall and F1-score, indicating its enhanced ability to correctly classify instances while maintaining robustness against misclassification.
@@
As shown in Fig. 9(b), our proposed method consistently outperforms MSNet and AMC_Net across all training set ratios (10%, 20%, 30%, and 100%). Notably, the model exhibits strong stability as the training data ratio decreases, maintaining high accuracy even under limited data conditions. Furthermore, Fig. 10(b) demonstrates that the model generalizes well to unseen data with 100 samples per class, highlighting the effective few-shot learning capability and reliable generalization of our model.
@@ IV-D3 Performance Comparison
Fig. 15 illustrates the performance of different spectrum sensing methods at various SNR levels. All methodologies exhibit poor performance when the SNR is between -10 dB and -4 dB. However, within the range of SNR from -10 dB to 0 dB, our proposed method demonstrates a higher accuracy rate. Beyond 0 dB, all methodologies achieve nearly perfect sensing accuracy.
Fig. 15 shows the receiver operating characteristic (ROC) curve of various spectrum sensing method at -4 dB SNR level. It highlights that our method achieves an AUC of 0.97 at -4 dB SNR, surpassing both LSTM and ResNet, demonstrating its superior accuracy in signal detection. At the same false alarm rate, our method also exhibits better detection performance, further underscoring its effectiveness in the SS task.
@@ IV-E2 Baseline
We leverage adversarial autoencoder (AAE) [aae] as the baseline model for AD task. The AAE is a generative model that integrates an autoencoder with a discriminator to enforce a specific feature distribution in the latent space. In AAE, the autoencoder consists of an encoder-decoder pair, where the encoder maps input data to a latent representation, and the decoder reconstructs the original input from this representation. Furthermore, a discriminator is employed to guide the encoder in shaping the latent space to follow a predefined distribution, such as a Gaussian or other structured priors.
@@ IV-E3 Performance Comparison
IV-F Joint Spectrum Sensing and Automatic Modulation Classification Task
SpectrumFM is further evaluated in a unified classification setting that combines spectrum sensing and automatic modulation classification [xing2024joint]. The joint task is formulated by augmenting the modulation label space with an additional class representing the idle state, so that the model directly predicts either “idle” or one of the active modulation types. This formulation allows spectrum occupancy and modulation type to be inferred simultaneously with a single classifier. As shown in Fig. 20, SpectrumFM achieves a competitive performance across various SNR levels, outperforming baseline methods in most regimes, particularly at moderate to high SNRs. The confusion matrix at 0 dB (Fig. 20) shows that SpectrumFM effectively distinguishes between the idle state and active modulation types, including OOK, 4ASK, and 8ASK, with low misclassification rates. These results demonstrate the feasibility and effectiveness of the unified classification framework for simultaneous spectrum sensing and modulation recognition.
IV-G Ablation Study
Figure 17: Accuracy of the joint spectrum sensing and automatic modulation classification task.
Figure 18: Confusion matrix of SpectrumFM for the joint task at 0 dB SNR.
Figure 19: Ablation study on the joint optimization of the masked reconstruction loss and the next-slot signal prediction loss.
Figure 20: Ablation study on the convolution module.
To better understand the contribution of each pretraining objective, we conduct an ablation study by separately evaluating the masked reconstruction loss and the next-slot signal prediction loss, in comparison with their joint optimization. The results are shown in Fig. 20.
When trained with only masked reconstruction loss, the model achieves a reasonable performance, reaching 0.6355 on AMC and 0.7961 on WTC, demonstrating the ability to capture local structural patterns and to enhance robustness against corrupted inputs. Training with only next-slot signal prediction loss yields 0.6274 on AMC and 0.8076 on WTC, showing that this objective effectively models sequential dependencies but does not consistently outperform masked reconstruction across different tasks. Most importantly, the combination of both objectives consistently provides the best results, indicating a clear complementary effect. Specifically, the masked reconstruction loss encourages the model to learn robust feature representations, while the next-slot signal prediction loss enforces the temporal consistency, and their synergy leads to stronger generalization and overall performance gains.
To further evaluate the architectural design, we also conduct an ablation experiment on the convolution module. As shown in Fig. 20, when this component is removed, the performance on downstream tasks drops notably, with AMC decreasing from 0.6372 to 0.6285 and WTC decreasing from 0.8253 to 0.7671. These results highlight the crucial role of convolutional layers in capturing fine-grained local spectral features, which complement the global dependencies modeled by MHSA and jointly contribute to the superior performance of SpectrumFM.
IV-H Discussion
@@
IV-I Limitations
Despite its strong performance, SpectrumFM has several limitations that need further investigation. First, the current pre-training dataset and model size, while sufficient for the demonstrated tasks, remain relatively limited, which constrains the model generalization ability and limits its performance in zero-shot scenarios. Second, although SpectrumFM has been evaluated on key downstream tasks such as AMC, WTC, SS, and AD, the range of supported tasks remains limited, and expanding the diversity of downstream applications could further validate its versatility. Third, the current model primarily operates on single-modality spectrum data. Extending SpectrumFM to support multi-modal inputs would enhance its applicability in more complex and heterogeneous wireless environments.
@@ V Conclusion
[Uncaptioned image] Fuhui Zhou (Senior Member, IEEE) is currently a Full Professor with Nanjing University of Aeronautics and Astronautics, Nanjing, China, where he is also with the Key Laboratory of Dynamic Cognitive System of Electromagnetic Spectrum Space. His research interests include cognitive radio, cognitive intelligence, knowledge graph, edge computing, and resource allocation. Prof. Zhou has published over 200 papers in internationally renowned journals and conferences in the field of communications. He has been selected for 1 ESI hot paper and 13 ESI highly cited papers. He has received 4 Best Paper Awards at international conferences such as IEEE Globecom and IEEE ICC. He was awarded as 2021 Most Cited Chinese Researchers by Elsevier, Stanford World’s Top 2% Scientists, IEEE ComSoc Asia-Pacific Outstanding Young Researcher and Young Elite Scientist Award of China and URSI GASS Young Scientist. He serves as an Editor of IEEE Wireless Communication Letters, IEEE Access and Physical Communications.
[Uncaptioned image] Chunyu Liu (Graduate Student Member, IEEE) is currently pursuing the Ph.D degree in electronic and information engineering with the School of Electronic and Information Engineering, Nanjing University of Aeronautics and Astronautics. His research interests include spectrum foundation models, knowledge graph, and wireless communication.
[Uncaptioned image] Hao Zhang (Member, IEEE) received the B.E. and M.Eng. from Nanchang University, China, in 2017 and 2020, respectively, and the Ph.D. degree from the College of Electronic and Information Engineering, Nanjing University of Aeronautics and Astronautics, Nanjing, China, in 2025. He is now a postdoctoral fellow with the College of Artificial Intelligence, Nanjing University of Aeronautics and Astronautics, Nanjing, China. He was a visiting Ph.D. student at the School of Electrical & Electronic Engineering, Nanyang Technological University, Singapore, in 2024. His research interests focus on deep learning, foundation models, wireless communication, wireless signal processing, and spectrum cognition.
[Uncaptioned image] Wei Wu (Member, IEEE) is currently an Full professor of the Nanjing University of Posts and Telecommunications. His research interests include spectrum sharing, semantic communication, knowledge graph, and physical layer security. He was the Technical Program Committee member for many international conferences, such as IEEE GLOBECOM and IEEE ICC. He is the editor of Physical Communications.
[Uncaptioned image] Qihui Wu (Fellow, IEEE) received the B.S. degree in communications engineering and the M.S. and Ph.D. degrees in communications and information systems from the Institute of Communications Engineering, Nanjing, China, in 1994, 1997, and 2000, respectively. Since May 2016, he has been a Full Professor with the College of Electronic and Information Engineering, Nanjing University of Aeronautics and Astronautics, Nanjing. From March 2011 to September 2011, he was an Advanced Visiting Scholar with the Stevens Institute of Technology, Hoboken, NJ, USA. His current research interests include the areas of wireless communications and statistical signal processing, with emphasis on system design of softwaredefined radio, cognitive radio, and smart radio.
[Uncaptioned image] Tony Q.S. Quek (Fellow, IEEE) received the B.E. and M.E. degrees in electrical and electronics engineering from the Tokyo Institute of Technology in 1998 and 2000, respectively, and the Ph.D. degree in electrical engineering and computer science from the Massachusetts Institute of Technology in 2008. Currently, he is the Cheng Tsang Man Chair Professor with Singapore University of Technology and Design (SUTD) and ST Engineering Distinguished Professor. He also serves as the Director of the Future Communications R&D Programme, the Head of ISTD Pillar, and the Deputy Director of the SUTD-ZJU IDEA. His current research topics include wireless communications and networking, network intelligence, non-terrestrial networks, open radio access network, and 6G. Dr. Quek has been actively involved in organizing and chairing sessions, and has served as a member of the Technical Program Committee as well as symposium chairs in a number of international conferences. He is currently serving as an Area Editor for the IEEE TRANSACTIONS ON WIRELESS COMMUNICATIONS. Dr. Quek was honored with the 2008 Philip Yeo Prize for Outstanding Achievement in Research, the 2012 IEEE William R. Bennett Prize, the 2015 SUTD Outstanding Education Awards - Excellence in Research, the 2020 IEEE Stephen O. Rice Prize, the 2020 Nokia Visiting Professor, and the 2022 IEEE Signal Processing Society Best Paper Award. He is a Fellow of IEEE and a Fellow of the Academy of Engineering Singapore.
[Uncaptioned image] Chan-Byoung Chae (Fellow, IEEE) received the Ph.D. degree in electrical and computer engineering from The University of Texas at Austin (UT), USA in 2008, where he was a member of wireless networking and communications group (WNCG). Prior to joining UT, he was a Research Engineer at the Telecommunications R&D Center, Samsung Electronics, Suwon, South Korea, from 2001 to 2005. He is currently an Underwood Distinguished Professor and Lee Youn Jae Fellow (Endowed Chair Professor) with the School of Integrated Technology, Yonsei University, South Korea. Before joining Yonsei, he was with Bell Labs, Alcatel-Lucent, Murray Hill, NJ, USA, from 2009 to 2011, as a Member of Technical Staff, and Harvard University, Cambridge, MA, USA, from 2008 to 2009, as a Post-Doc., Fellow and Lecturer. Dr. Chae was a recipient/co-recipient of the Korean Ministry of ICT and Science Award in 2024, the Korean Ministry of Education Award in 2024, the KICS Haedong Scholar Award in 2023, the CES Innovation Award in 2023, the IEEE ICC Best Demo Award in 2022, the IEEE WCNC Best Demo Award in 2020, the Best Young Engineer Award from the National Academy of Engineering of Korea (NAEK) in 2019, the IEEE DySPAN Best Demo Award in 2018, the IEEE/KICS Journal of Communications and Networks Best Paper Award in 2018, the IEEE INFOCOM Best Demo Award in 2015, the IEIE/IEEE Joint Award for Young IT Engineer of the Year in 2014, the KICS Haedong Young Scholar Award in 2013, the IEEE Signal Processing Magazine Best Paper Award in 2013, the IEEE ComSoc AP Outstanding Young Researcher Award in 2012, and the IEEE VTS Dan. E. Noble Fellowship Award in 2008. Dr. Chae has held several editorial positions, including Editor-in-Chief of the IEEE TRANSACTIONS ON MOLECULAR, BIOLOGICAL, AND MULTISCALE COMMUNICATIONS, Senior Editor of the IEEE WIRELESS COMMUNICATIONS LETTERS, and Editor of the IEEE COMMUNICATIONS MAGAZINE, IEEE TRANSACTIONS ON WIRELESS COMMUNICATIONS, and IEEE WIRELESS COMMUNICATIONS LETTERS. He was an IEEE ComSoc Distinguished Lecturer from 2020 to 2023 and is an IEEE VTS Distinguished Lecturer from 2024 to 2025. He is an elected member of the National Academy of Engineering of Korea.